# Phase transition and universality of the majority-rule model on complex networks

### 21 Feb 2024 v1

@@ Abstract
We investigate the phenomena of order-disorder phase transition and the universality of the majority rule model defined on three complex networks, namely the Barabási-Albert, Watts–Strogatz, and Erdős-Rényi networks. Assume each agent holds two possible opinions distributed randomly across the networks’ nodes. Agents adopt anticonformity and independence behaviors, represented by the probability $p$, where with a probability $p$, agents adopt anticonformity or independence behavior. Based on our numerical simulation results and finite-size scaling analysis, it is found that the model undergoes a continuous phase transition for all networks, with critical points for the independence model greater than those for the anticonformity model in all three networks. We obtain critical exponents identical to the opinion dynamics model defined on a complete graph, indicating that the model exhibits the same universality class as the mean-field Ising model.
@@ 1 Introduction
One approach to understanding socio-political phenomena based on physics principles is proposing an opinion dynamics model. This approach leverages statistical physics principles to analyze and interpret the complex interactions within socio-political systems galam1982sociophysics ; galam1999application ; castellano2009statistical ; galam2012socio ; sen2014sociophysics ; galam2017trump ; schweitzer2018sociophysics . In opinion dynamics models, the micro-level interactions between agents or individuals resemble spin interactions observed in the Ising model. Moreover, the opinion dynamics models are characterized by rules that dictate the interactions among individuals to reach a particular outcome, such as achieving a consensus state. Models of opinion dynamics represent a topic of considerable interest within sociophysics. These models offer insightful perspectives on how individual opinions evolve and interact within social networks, contributing to our understanding of collective behavior in various sociological contexts. For example, the classic voter model illustrates the opinion formation process in a simplified setting. In this model, each agent possesses two potential opinions, influenced predominantly by the opinions of their immediate neighbors. Despite its simplicity, the voter model can exhibit complex behavior. Over time, the system may reach a consensus (all agents sharing the same opinion) or remain in a state of coexistence where different opinions persist indefinitely, depending on the structure of the underlying network and initial conditions holley1975ergodic . The Axelrod model represents another sociophysical framework, elucidating the dissemination of cultural traits through local interactions. This model underscores the emergence of distinct cultural regions grounded in the principle of homophily. It demonstrates the formation and evolution of cultural domains, where the emergence of culturally similar groups leads to the development of clusters or regions characterized by shared cultural traits axelrod1997dissemination . Two other popular opinion dynamics models are the Sznajd and majority-rule models. These models offer distinct perspectives on how opinions form and evolve within social groups. The Sznajd Model provides an insightful representation of how pairs of like-minded individuals influence their neighbors, leading to a range of macroscopic societal states, including consensus, polarization, or fragmentation sznajd2000opinion . The majority-rule model is constructed to replicate the dynamics through which collective opinions or decisions materialize within a group, predominantly influenced by the majority’s preference galam1986majority ; krapivsky2003dynamics . Various discrete and continuous opinion models have been proposed and developed to understand social aspects thoroughly. Discrete opinion models, such as the Sznajd model sznajd2000opinion , the $q$-voter model castellano2009nonlinear , and the Biswas-Sen model biswas2009model , typically characterize the opinions of individuals (agents) as being represented by a limited number of distinct choices. In contrast, continuous opinion models, such as the DeGroot model degroot1974reaching and the Deffuant–Weisbuch model deffuant2000mixing , consider opinions as part of a continuous spectrum. Each opinion dynamics model framework possesses its advantages, depending on the social context under consideration. For instance, discrete models are more suitable for applications like voting or surveys concerning the selection of specific candidates, while continuous models are more apt for use in social psychological and sentiment analysis, where opinions are more complex and gradational. In addition to the nature of opinions, the network structure governing agent interactions plays a pivotal role. Individuals maintain diverse social connections in real-world social contexts, making conceptualizing opinion dynamics models within heterogeneous or complex networks more pragmatic. These include networks such as the Barabási-Albert (B-A) albert2002statistical , Watts-Strogatz (W-S) watts1998collective , and Erdős-Rényi (E-R) models erdHos1960evolution , as opposed to homogeneous networks typified by regular lattices and complete graphs. Such an approach aligns more closely with the intricate and multifaceted nature of real-world social networks newman2018networks . From a physics perspective, several interesting physical features exist to explore, such as phase transition phenomena and universality cardy1996scaling ; crokidakis2015inflexibility ; mukherjee2016disorder ; oestereich2019three ; calvelli2019phase ; muslim2020phase ; muslim2021phase ; pires2022double ; muslim2022opinion ; schawe2022higher . The investigation of the phase transition phenomenon is compelling due to its capacity to elucidate shifts in social dynamics, such as consensus-status quo or agreement-disagreement, manifested in continuous and discontinuous phase transitions. These features have been extensively studied in opinion dynamics models with various social scenarios and network topologies. Examples include the examination of universality in the Sznajd model on complete graphs with two distinct configurations muslim2020phase , the analysis of the q-voter model using a pair approximation approach on multiplex networks gradowski2020pair , the study of phase transition phenomena due to anticonformity and independence behaviors on regular lattices mulya2023destructive , the investigation of phase transition phenomena in Barabási-Albert networks based on the Biswas–Chatterjee–Sen Model alencar2023opinion , and the analysis of the Galam model with contrarian on complete graphs under the influence of time-varying external fields gimenez2023contrarian . They also help to understand macroscopically the dynamics that occur, even if the microscopic interactions among individuals can be intricate and complex. Another interesting opinion dynamics model is the majority rule model, mainly when defined on complex networks. Phenomena of both continuous and discontinuous phase transitions from order to disorder have been observed in random graphs pereira2005majority ; lima2008majority , small-world networks campos2003small ; luz2007majority ; stone2015majority ; zubillaga2022three , and scale-free networks chen2015critical ; vilela2020three ; alencar2023droplet . These studies collectively reveal a phase transition phenomenon at a specific critical point. However, to the best of our knowledge, there has been no comparative study exploring the impact of two types of noise defined within three networks simultaneously: B-A network (scale-free), W-S (small-world), and E-R (random graph) and how the universality class of the model manifests across these three networks, which is the main study in this paper. In a social context, both types of noise are correlated with social behaviors that disrupt consensus, where consensus decreases as increases the social behaviors, leading to a social phase changes at a specific threshold nyczka2013anticonformity ; nowak2019homogeneous . By borrowing terms from social psychology, these disruptive social behaviors are anticonformity and independence. Both behaviors are fundamentally contrary to conformity behavior, which always adopts the majority opinion. The difference is that anticonformity behavior evaluates the surrounding situation and acts oppositely, while independent behavior acts independently without any pressure from any party willis1963two ; willis1965conformity ; macdonald2004expanding ; nail2011development . The meanings of these two behaviors can minimize the complete consensus state where all agents eventually have the same unanimous opinion. Various scenarios of opinion dynamics models with anticonformity and independence behaviors and the impact of both behaviors can be found in Refs. sznajd2011phase ; nyczka2013anticonformity ; javarone2014social ; crokidakis2015inflexibility ; chmiel2015phase ; abramiuk2019independence ; muslim2020phase ; nowak2021discontinuous ; civitarese2021external ; muslim2021phase ; muslim2022opinion ; muslim4241509phase . For instance, in Ref. sznajd2011phase , the authors observed the phenomenon of continuous phase transition within the Sznajd model with independence behavior defined on a complete graph, 1D lattice, and 2D lattice. Continuous phase transition phenomena and universality classes were observed in the majority-rule model with nonconformity behavior on a complete graph, body-centered cubic (bcc) lattice, and triangular lattice crokidakis2015inflexibility . Our recent studies also investigated the continuous phase transition phenomenon and universality classes of the Sznajd model with anticonformity and independence on a complete graph with two scenarios muslim2022opinion and the majority-rule model with anticonformity and independence behaviors defined on a 2D lattice with square formation muslim4241509phase . In general, the behaviors of anticonformity and independence in the mentioned models are represented by the probability parameter $p$, where agents adopt anticonformity or independence behaviors with a probability of $p$. Order-disorder phase transitions within opinion dynamics models are sometimes interpreted to understand further social phase changes, such as consensus-status quo and agreement-disagreement, as observed in several studies above. As mentioned earlier, understanding such social phase changes would be more intriguing if opinion dynamics models were defined on complex networks that can represent real social networks. This paper investigates destructive social behaviors that disrupt homogeneity, namely anticonformity and independence behaviors. Both behaviors can be considered as social noise which works similarly to temperature in the Ising model. These behaviors are represented by the probability $p$, where with a probability $p$, agents adopt anticonformity or independence behavior. Our preliminary results for this model on homogeneous networks such as 2D lattice, 3D lattice, and a complete graph are discussed in Ref. mulya2023destructive . Our numerical simulation results show that the critical point causing the model to undergo a continuous phase transition is smaller for the model with anticonformity than for the model with independence. From a socio-political perspective, anti-conformist individuals are likelier to make the community more chaotic than independent individuals. Additionally, one can argue that the majority-minority opinions exist below the critical point, and at the critical point $p_{c}$, the system is in a stalemate situation or status quo. Based on finite-size scaling analysis, we obtained best-fit critical exponents that caused the collapse of all data. The critical exponents indicate that the model with anticonformity and independence on all three networks is identical and belongs to the same universality class as the opinion dynamics model defined on a complete graph and the mean-field Ising model.
@@ 2 Model and Method
The majority rule model is an opinion dynamics model that assumes the majority opinion always prevails. This model makes sense when considering a group of individuals with an odd number of members interacting with each other. Suppose we assume that only the majority rule mechanism will always win in every interaction. In that case, naturally, a homogeneous state will always occur, and the final state depends on the initial state of the system. For instance, if we examine two competing opinions (opinions A and B) within a population, the opinion with the initially larger size will ultimately prevail. In addition to depending on the initial population size and system size, external parameters, such as the influence of an external field (mass media), play a significant role in influencing the consensus state of the system sirbu2017opinion ; li2020effect ; azhari2023external ; muslim2024mass . Typically, in the absence of external influence and with a sufficiently large population size, it is observed that at an initial population fraction of $c=0.5$, the system possesses an equal probability of evolving toward a complete consensus of either $m=+1$ (all agents have the same state or opinion up) or $m=-1$ (all agents have the same state or opinion down). The point $c=0.5$ can be called the ‘critical point’ or separator point between the two consensus states muslim2024mass . Social interpretations, such as changes in consensus-status quo or vice versa, can be analyzed through changes in the order parameter (magnetization) concerning the noise parameter that disrupts regularity. This noise parameter functions similarly to temperature in the Ising model; at low temperatures, the dominant factor is the binding energy allowing all spins to align in one direction (ferromagnetic), but at high temperatures, thermal fluctuations disrupt this order, causing the material to lose its magnetization newell1953theory . To obtain a more relevant understanding of social phenomena, the opinion dynamics model is more realistic when analyzed on complex networks rather than homogeneous networks such as regular lattices or complete graphs. The reality is that the structure of complex networks is much more realistic than homogeneous networks, where each node has varying numbers of nearest neighbors (depicting different social relationships). As mentioned earlier, statistical physics features such as phase transitions make more sense when translated into socio-political language. In this paper, we study the order-disorder phase transition of the majority-rule model with anticonformity and independence on three complex network models, namely B-A, W-S, and E-R networks. The mean degree for the B-A, W-S, and E-R networks are $\langle k\rangle\approx 10,\langle k\rangle\approx 4$, and $\langle k\rangle=r(N-1)$ respectively, where $r=0.1$, ensuring that each node has at least two directly connected neighboring nodes for all populations $N$ utilized. We consider three agents interacting with each other, and these agents also adopt destructive social behaviors, namely anticonformity and independence. The model in this paper can be described as follows:
@@
    Three agents, for example, $S_{0},S_{1}$, and $S_{2}$, are randomly selected to follow the majority rule, as illustrated in Fig. 1.
@@
Figure 1: Illustration of the selection of three agents at random (initial and updated state): agent $S_{0}$ is initially chosen, followed by the random selection of its two neighboring agents, $S_{1}$ and $S_{2}$ (solid edges), who interact based on the majority rule. The red and orange colors represent agents with opinions or states of “up” and “down,” respectively. All agents are distributed randomly across the networks’ nodes.
@@ 3 Result and Discussion
We have presented preliminary simulation results for the model in this paper, demonstrating that the model undergoes a continuous phase transition for all networks, including B-A, W-S, and E-R networks, for both the models with anticonformist and independent agents mulya2023destructive . In that article, we did not estimate the critical points and critical exponent parameters on the three networks, which are the main focus of discussion in this paper. All results presented in this paper are outcomes of numerical simulations. We conducted numerical simulations with various population sizes $N$ to estimate critical points and exponents using Eqs. (2) through (5). The order parameter $m$ was computed by Eq. (1), while susceptibility $\chi$ and Binder cumulant $U$ were determined using Eqs. (6) and (7) respectively. Eqs. (2) through (4) can be employed to obtain critical exponent values $\nu,~{}\beta,$ and $\gamma$. This can be achieved by utilizing the critical point value $p_{c}$ obtained from the intersection of lines of the Binder cumulant $U$ versus probability $p$ and calculating the pseudo-critical point $p_{c}(N)$ for each population size $N$ based on the maximum susceptibility $\chi$ (see Fig. 2). Furthermore, the order parameter $m$ at each population’s pseudo-critical point $p_{c}(N)$ was also obtained. To determine the ratios $1/\nu,~{}\gamma/\nu,$ and $\beta/\nu$, ordinary linear regression was employed to extract gradient values from the data. This approach simplifies the process of obtaining these critical exponent ratios.
@@
In general, we computed the average for each data point corresponding to $m,\chi,$ and $U$ over more than $10^{6}$ independent realizations to ensure the accuracy of our results. Each data point was calculated when the simulation had reached its steady state. Our numerical simulation results for scaling plot of $U,m,$ and $\chi$ on the B-A network are exhibited in Fig. 4. Panels (a) - (c) correspond to the model with anticonformist agents, while panels (d) - (f) correspond to the model with independent agents. The critical point $p_{c}$ is determined from the intersection of lines of the Binder cumulant $U$ versus probability $p$, yielding $p_{c}=0.203(3)$ for the anticonformity model and $p_{c}=0.301(2)$ for the independence model. These data indicate that the models undergo a continuous phase transition at the mentioned critical points. These data also lead to the collapse of all data for different population sizes $N$ around the critical point $p_{c}$. A similar approach was applied to the model with anticonformist agents on the B-A network. We obtained the critical exponents as follows: $1/\nu=0.495(3)\rightarrow\nu=2.020(3),\gamma/\nu=0.503(2)\rightarrow\gamma=1.016(2)$ and $\beta/\nu=0.252(3)\rightarrow\beta=0.509(3)$. The model on W-S and E-R networks also undergoes a continuous phase transition with critical points as follows: For the model on the W-S network with anticonformist and independent agents, the critical points are $p_{c}=0.116(2)$ and $p_{c}=0.203(2)$, respectively. For the model on the E-R network with anticonformist and independent agents, the critical points are $p_{c}=0.255(3)$ and $p_{c}=0.337(3)$, respectively. Interestingly, based on the finite-size scaling analysis, the models on the W-S and E-R networks exhibit similar critical exponents to the model on the B-A network. In detail, the critical exponents for the model with anticonformist agents on the W-S network are $1/\nu=0.493(5)\rightarrow\nu=2.028(5)$, $\gamma/\nu=0.494(4)\rightarrow\gamma=1.001(4)$, and $\beta/\nu=0.247(5)\rightarrow\beta=0.501(5)$ (see panels (a) - (c) of Fig. 5 for the scaling plot). Furthermore, the critical exponents for the model with independent agents on the W-S network are $1/\nu=0.502(4)\rightarrow\nu=1.992(4)$, $\gamma/\nu=0.501(2)\rightarrow\gamma=0.998(2)$, and $\beta/\nu=0.253(3)\rightarrow\beta=0.503(3)$ (see panels (d) - (f) of Fig. 5 for the scaling plot). For the models on the E-R network, we obtain critical exponents for the model with anticonformist agents as follows: $1/\nu=0.497(3)\rightarrow\nu=2.012(3)$, $\gamma/\nu=0.493(2)\rightarrow\gamma=0.991(2)$, and $\beta/\nu=0.248(4)\rightarrow\beta=0.498(4)$ (see panels (a) - (c) of Fig. 6 for the scaling plot). Meanwhile, for the model with independent agents on the E-R network, its critical exponents are $1/\nu=0.498(3)\rightarrow\nu=2.000(3)$, $\gamma/\nu=0.496(2)\rightarrow\gamma=0.995(2)$, and $\beta/\nu=0.251(3)\rightarrow\beta=0.502(3)$ (see panels (d) - (f) of Fig. 6 for the scaling plot).

### 26 Feb 2024 v2

@@ Abstract
We investigate the phenomena of order-disorder phase transition and the universality of the majority-rule model defined on three complex networks, namely the Barabási-Albert, Watts–Strogatz, and Erdős-Rényi networks. Assume each agent holds two possible opinions distributed randomly across the networks’ nodes. Agents adopt anticonformity and independence behaviors, represented by the probability $p$, where with a probability $p$, agents adopt anticonformity or independence behavior. Based on our numerical simulation results and finite-size scaling analysis, it is found that the model undergoes a continuous phase transition for all networks, with critical points for the independence model greater than those for the anticonformity model in all three networks. We obtain critical exponents identical to the opinion dynamics model defined on a complete graph, indicating that the model exhibits the same universality class as the mean-field Ising model.
@@
[brin] organization=Research Center for Quantum Physics, National Research and Innovation Agency (BRIN), city=South Tangerang, postcode=15314, country=Indonesia
@@
[uty] organization=Department of Industrial Engineering, University of Technology Yogyakarta, city=Yogyakarta, postcode=55285, country=Indonesia
@@ 1 Introduction
One approach to understanding socio-political phenomena based on physics principles is proposing an opinion dynamics model. This approach leverages statistical physics principles to analyze and interpret the complex interactions within socio-political systems galam1982sociophysics ; galam1999application ; castellano2009statistical ; galam2012socio ; sen2014sociophysics ; galam2017trump ; schweitzer2018sociophysics . In opinion dynamics models, the micro-level interactions between agents or individuals resemble spin interactions observed in the Ising model. Moreover, the opinion dynamics models are characterized by rules that dictate the interactions among individuals to reach a particular outcome, such as achieving a consensus state.
Models of opinion dynamics represent a topic of considerable interest within sociophysics. These models offer insightful perspectives on how individual opinions evolve and interact within social networks, contributing to our understanding of collective behavior in various sociological contexts. For example, the classic voter model illustrates the opinion formation process in a simplified setting. In this model, each agent possesses two potential opinions, influenced predominantly by the opinions of their immediate neighbors. Despite its simplicity, the voter model can exhibit complex behavior. Over time, the system may reach a consensus (all agents sharing the same opinion) or remain in a state of coexistence where different opinions persist indefinitely, depending on the structure of the underlying network and initial conditions holley1975ergodic . The Axelrod model represents another sociophysical framework, elucidating the dissemination of cultural traits through local interactions. This model underscores the emergence of distinct cultural regions grounded in the principle of homophily. It demonstrates the formation and evolution of cultural domains, where the emergence of culturally similar groups leads to the development of clusters or regions characterized by shared cultural traits axelrod1997dissemination . Two other popular models of opinion dynamics are the Sznajd model and the majority-rule model, also referred to as the Galam majority-rule model or the Galam model of opinion dynamism. These models offer distinct perspectives on how opinions form and evolve within social groups. The Sznajd Model provides an insightful representation of how pairs of like-minded individuals influence their neighbors, leading to a range of macroscopic societal states, including consensus, polarization, or fragmentation sznajd2000opinion . The majority-rule model is constructed to replicate the dynamics through which collective opinions or decisions materialize within a group, predominantly influenced by the majority’s preference galam1986majority ; krapivsky2003dynamics .
Various discrete and continuous opinion models have been proposed and developed to understand social aspects thoroughly. Discrete opinion models, such as the Sznajd model sznajd2000opinion , the $q$-voter model castellano2009nonlinear , and the Biswas-Sen model biswas2009model , typically characterize the opinions of individuals (agents) as being represented by a limited number of distinct choices. In contrast, continuous opinion models, such as the DeGroot model degroot1974reaching and the Deffuant–Weisbuch model deffuant2000mixing , consider opinions as part of a continuous spectrum. Each opinion dynamics model framework possesses its advantages, depending on the social context under consideration. For instance, discrete models are more suitable for applications like voting or surveys concerning the selection of specific candidates, while continuous models are more apt for use in social psychological and sentiment analysis, where opinions are more complex and gradational.
In addition to the nature of opinions, the network structure governing agent interactions plays a pivotal role. Individuals maintain diverse social connections in real-world social contexts, making conceptualizing opinion dynamics models within heterogeneous or complex networks more pragmatic. These include networks such as the Barabási-Albert (B-A) albert2002statistical , Watts-Strogatz (W-S) watts1998collective , and Erdős-Rényi (E-R) models erdHos1960evolution , as opposed to homogeneous networks typified by regular lattices and complete graphs. Such an approach aligns more closely with the intricate and multifaceted nature of real-world social networks newman2018networks .
From a physics perspective, several interesting physical features exist to explore, such as phase transition phenomena and universality cardy1996scaling ; crokidakis2015inflexibility ; mukherjee2016disorder ; oestereich2019three ; calvelli2019phase ; muslim2020phase ; muslim2021phase ; pires2022double ; muslim2022opinion ; schawe2022higher . The investigation of the phase transition phenomenon is compelling due to its capacity to elucidate shifts in social dynamics, such as consensus-status quo or agreement-disagreement, manifested in continuous and discontinuous phase transitions. These features have been extensively studied in opinion dynamics models with various social scenarios and network topologies. Examples include the examination of universality in the Sznajd model on complete graphs with two distinct configurations muslim2020phase , the analysis of the q-voter model using a pair approximation approach on multiplex networks gradowski2020pair , the study of phase transition phenomena due to anticonformity and independence behaviors on regular lattices mulya2023destructive , the investigation of phase transition phenomena in Barabási-Albert networks based on the Biswas–Chatterjee–Sen Model alencar2023opinion , and the analysis of the Galam model with contrarian on complete graphs under the influence of time-varying external fields gimenez2023contrarian . They also help to understand macroscopically the dynamics that occur, even if the microscopic interactions among individuals can be intricate and complex.
Another interesting opinion dynamics model is the majority-rule model, mainly when defined on complex networks. Phenomena of both continuous and discontinuous phase transitions from order to disorder have been observed in random graphs pereira2005majority ; lima2008majority , small-world networks campos2003small ; luz2007majority ; stone2015majority ; zubillaga2022three , and scale-free networks chen2015critical ; vilela2020three ; alencar2023droplet . These studies collectively reveal a phase transition phenomenon at a specific critical point. However, to the best of our knowledge, there has been no comparative study exploring the impact of two types of noise defined within three networks simultaneously: B-A network (scale-free), W-S (small-world), and E-R (random graph) and how the universality class of the model manifests across these three networks, which is the main study in this paper. In a social context, both types of noise are correlated with social behaviors that disrupt consensus, where consensus decreases as increases the social behaviors, leading to a social phase changes at a specific threshold nyczka2013anticonformity ; nowak2019homogeneous .
By borrowing terms from social psychology, these disruptive social behaviors are anticonformity and independence. Both behaviors are fundamentally contrary to conformity behavior, which always adopts the majority opinion. The difference is that anticonformity behavior evaluates the surrounding situation and acts oppositely, while independent behavior acts independently without any pressure from any party willis1963two ; willis1965conformity ; macdonald2004expanding ; nail2011development . The meanings of these two behaviors can minimize the complete consensus state where all agents eventually have the same unanimous opinion. Various scenarios of opinion dynamics models with anticonformity and independence behaviors and the impact of both behaviors can be found in Refs. sznajd2011phase ; nyczka2013anticonformity ; javarone2014social ; crokidakis2015inflexibility ; chmiel2015phase ; abramiuk2019independence ; muslim2020phase ; nowak2021discontinuous ; civitarese2021external ; muslim2021phase ; muslim2022opinion ; muslim4241509phase . For instance, in Ref. sznajd2011phase , the authors observed the phenomenon of continuous phase transition within the Sznajd model with independence behavior defined on a complete graph, 1D lattice, and 2D lattice. Continuous phase transition phenomena and universality classes were observed in the majority-rule model with nonconformity behavior on a complete graph, body-centered cubic (bcc) lattice, and triangular lattice crokidakis2015inflexibility . Our recent studies also investigated the continuous phase transition phenomenon and universality classes of the Sznajd model with anticonformity and independence on a complete graph with two scenarios muslim2022opinion and the majority-rule model with anticonformity and independence behaviors defined on a 2D lattice with square formation muslim4241509phase . In general, the behaviors of anticonformity and independence in the mentioned models are represented by the probability parameter $p$, where agents adopt anticonformity or independence behaviors with a probability of $p$. Order-disorder phase transitions within opinion dynamics models are sometimes interpreted to understand further social phase changes, such as consensus-status quo and agreement-disagreement, as observed in several studies above.
As mentioned earlier, understanding such social phase changes would be more intriguing if opinion dynamics models were defined on complex networks that can represent real social networks. This paper investigates destructive social behaviors that disrupt homogeneity, namely anticonformity and independence behaviors. Both behaviors can be considered as social noise which works similarly to temperature in the Ising model. These behaviors are represented by the probability $p$, where with a probability $p$, agents adopt anticonformity or independence behavior. Our preliminary results for this model on homogeneous networks such as 2D lattice, 3D lattice, and a complete graph are discussed in Ref. mulya2023destructive .
Our numerical simulation results show that the critical point causing the model to undergo a continuous phase transition is smaller for the model with anticonformity than for the model with independence. From a socio-political perspective, anti-conformist individuals are likelier to make the community more chaotic than independent individuals. Additionally, one can argue that the majority-minority opinions exist below the critical point, and at the critical point $p_{c}$, the system is in a stalemate situation or status quo. Based on finite-size scaling analysis, we obtained best-fit critical exponents that caused the collapse of all data. The critical exponents indicate that the model with anticonformity and independence on all three networks is identical and belongs to the same universality class as the opinion dynamics model defined on a complete graph and the mean-field Ising model.
@@ 2 Model and Method
The majority-rule model is an opinion dynamics model that assumes the majority opinion always prevails. This model makes sense when considering a group of individuals with an odd number of members interacting with each other. Suppose we assume that only the majority-rule mechanism will always win in every interaction. In that case, naturally, a homogeneous state will always occur, and the final state depends on the initial state of the system. For instance, if we examine two competing opinions (opinions A and B) within a population, the opinion with the initially larger size will ultimately prevail.
@@
In addition to depending on the initial population size and system size, external parameters, such as the influence of an external field (mass media), play a significant role in influencing the consensus state of the system sirbu2017opinion ; li2020effect ; azhari2023external ; muslim2024mass . Typically, in the absence of external influence and with a sufficiently large population size, it is observed that at an initial population fraction of $c=0.5$, the system possesses an equal probability of evolving toward a complete consensus of either $m=+1$ (all agents have the same state or opinion up) or $m=-1$ (all agents have the same state or opinion down). The point $c=0.5$ can be called the ‘critical point’ or separator point between the two consensus states muslim2024mass . Social interpretations, such as changes in consensus-status quo or vice versa, can be analyzed through changes in the order parameter (magnetization) concerning the noise parameter that disrupts regularity. This noise parameter functions similarly to temperature in the Ising model; at low temperatures, the dominant factor is the binding energy allowing all spins to align in one direction (ferromagnetic), but at high temperatures, thermal fluctuations disrupt this order, causing the material to lose its magnetization newell1953theory .
To obtain a more relevant understanding of social phenomena, the opinion dynamics model is more realistic when analyzed on complex networks rather than homogeneous networks such as regular lattices or complete graphs. The reality is that the structure of complex networks is much more realistic than homogeneous networks, where each node has varying numbers of nearest neighbors (depicting different social relationships). As mentioned earlier, statistical physics features such as phase transitions make more sense when translated into socio-political language.
In this paper, we study the order-disorder phase transition of the majority-rule model with anticonformity and independence on three complex network models, namely B-A, W-S, and E-R networks. The mean degree for the B-A, W-S, and E-R networks are $\langle k\rangle\approx 10,\langle k\rangle\approx 4$, and $\langle k\rangle=r(N-1)$ respectively, where $r=0.1$, ensuring that each node has at least two directly connected neighboring nodes for all populations $N$ utilized. We consider three agents interacting with each other, and these agents also adopt destructive social behaviors, namely anticonformity and independence.
The model in this paper can be described as follows:
@@
    Three agents, for example, $S_{0},S_{1}$, and $S_{2}$, are randomly selected to follow the majority-rule, as illustrated in Fig. 1.
@@
Figure 1: Illustration of the selection of three agents at random (initial and updated state): agent $S_{0}$ is initially chosen, followed by the random selection of its two neighboring agents, $S_{1}$ and $S_{2}$ (solid edges), who interact based on the majority-rule. The red and orange colors represent agents with opinions or states of “up” and “down,” respectively. All agents are distributed randomly across the networks’ nodes.
@@ 3 Result and Discussion
We have presented preliminary simulation results for the model in this paper, demonstrating that the model undergoes a continuous phase transition for all networks, including B-A, W-S, and E-R networks, for both the models with anticonformist and independent agents mulya2023destructive . In that article, we did not estimate the critical points and critical exponent parameters on the three networks, which are the main focus of discussion in this paper. All results presented in this paper are outcomes of numerical simulations.
We conducted numerical simulations with various population sizes $N$ to estimate critical points and exponents using Eqs. (2) through (5). The order parameter $m$ was computed by Eq. (1), while susceptibility $\chi$ and Binder cumulant $U$ were determined using Eqs. (6) and (7) respectively. Eqs. (2) through (4) can be employed to obtain critical exponent values $\nu,~{}\beta,$ and $\gamma$. This can be achieved by utilizing the critical point value $p_{c}$ obtained from the intersection of lines of the Binder cumulant $U$ versus probability $p$ and calculating the pseudo-critical point $p_{c}(N)$ for each population size $N$ based on the maximum susceptibility $\chi$ (see Fig. 2). Furthermore, the order parameter $m$ at each population’s pseudo-critical point $p_{c}(N)$ was also obtained. To determine the ratios $1/\nu,~{}\gamma/\nu,$ and $\beta/\nu$, ordinary linear regression was employed to extract gradient values from the data. This approach simplifies the process of obtaining these critical exponent ratios.
@@
In general, we computed the average for each data point corresponding to $m,\chi,$ and $U$ over more than $10^{6}$ independent realizations to ensure the accuracy of our results. Each data point was calculated when the simulation had reached its steady state. Our numerical simulation results for scaling plot of $U,m,$ and $\chi$ on the B-A network are exhibited in Fig. 4. Panels (a) - (c) correspond to the model with anticonformist agents, while panels (d) - (f) correspond to the model with independent agents. The critical point $p_{c}$ is determined from the intersection of lines of the Binder cumulant $U$ versus probability $p$, yielding $p_{c}=0.203(3)$ for the anticonformity model and $p_{c}=0.301(2)$ for the independence model. These data indicate that the models undergo a continuous phase transition at the mentioned critical points. These data also lead to the collapse of all data for different population sizes $N$ around the critical point $p_{c}$. A similar approach was applied to the model with anticonformist agents on the B-A network. We obtained the critical exponents as follows: $1/\nu=0.495(3)\rightarrow\nu=2.020(3),\gamma/\nu=0.503(2)\rightarrow\gamma=1.016(2)$ and $\beta/\nu=0.252(3)\rightarrow\beta=0.509(3)$.
The model on W-S and E-R networks also undergoes a continuous phase transition with critical points as follows: For the model on the W-S network with anticonformist and independent agents, the critical points are $p_{c}=0.116(2)$ and $p_{c}=0.203(2)$, respectively. For the model on the E-R network with anticonformist and independent agents, the critical points are $p_{c}=0.255(3)$ and $p_{c}=0.337(3)$, respectively. Interestingly, based on the finite-size scaling analysis, the models on the W-S and E-R networks exhibit similar critical exponents to the model on the B-A network. In detail, the critical exponents for the model with anticonformist agents on the W-S network are $1/\nu=0.493(5)\rightarrow\nu=2.028(5)$, $\gamma/\nu=0.494(4)\rightarrow\gamma=1.001(4)$, and $\beta/\nu=0.247(5)\rightarrow\beta=0.501(5)$ (see panels (a) - (c) of Fig. 5 for the scaling plot). Furthermore, the critical exponents for the model with independent agents on the W-S network are $1/\nu=0.502(4)\rightarrow\nu=1.992(4)$, $\gamma/\nu=0.501(2)\rightarrow\gamma=0.998(2)$, and $\beta/\nu=0.253(3)\rightarrow\beta=0.503(3)$ (see panels (d) - (f) of Fig. 5 for the scaling plot). For the models on the E-R network, we obtain critical exponents for the model with anticonformist agents as follows: $1/\nu=0.497(3)\rightarrow\nu=2.012(3)$, $\gamma/\nu=0.493(2)\rightarrow\gamma=0.991(2)$, and $\beta/\nu=0.248(4)\rightarrow\beta=0.498(4)$ (see panels (a) - (c) of Fig. 6 for the scaling plot). Meanwhile, for the model with independent agents on the E-R network, its critical exponents are $1/\nu=0.498(3)\rightarrow\nu=2.000(3)$, $\gamma/\nu=0.496(2)\rightarrow\gamma=0.995(2)$, and $\beta/\nu=0.251(3)\rightarrow\beta=0.502(3)$ (see panels (d) - (f) of Fig. 6 for the scaling plot).